Which of the following are valid arithmetic statements?

PLEASE EXPLAIN WHY

a. x = 4 * 5

b. ansA && ansB = true

c. 2 + 10 / 2 = c

d. all of the above
Answer #1
c. 2 + 10 / 2 = c
This is not a valid arithmetic statement.
Because, value should be right side of the assignment ( = ) operator and variable should at the left side as shown below.
c = 2 + 10 / 2;

b. ansA && ansB = true
we have to use == to compare two values instead of =
Because, = is an assignment operator where as == is a comparative operator.

So, Option a is valid arithmetic statement.

a. x = 4 * 5